OpenStudy (anonymous):

The area of a rectangle is 48 square feet. The length is two feet longer than the width. What is the width of the rectangle?

OpenStudy (armyrangers):

What are the factors of 48

OpenStudy (anonymous):

none were given

OpenStudy (armyrangers):

You should know the factors of 48

OpenStudy (anonymous):

1,2,3,4,6,8,12,16,24,48

OpenStudy (anonymous):

Well, my way is a little weird but here's how I'd solve it: Set up an algebraic equation: We know that area=width*length We also know that the area is 48, and that the length is 2 feet longer than the width. So we now have: 48 =N*(N+2) Because the one side is 2 feet longer than the other, and we don't know the length of either. 48=N^2 +2N N=6. So the width is 6, and the length is 8 (6+2)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I can see what army was trying to show you. You could see that 6 and 8 were both factors of 48, and that 8 was 2 bigger than 6, so your length was 8 and your width was 6. There explanation was a little better. :)

OpenStudy (armyrangers):

Thanks fort he pat on the back